# Hysteretic Excitation in Non-collinear Antiferromagnetic Spin-Torque Oscillators: A Terminal Velocity Motion Perspective

### v3

@@ II.6 Threshold and Critical Current Densities Verified by the TVM/Macrospin Simulations
Furthermore, the model comparison provides insight into the oscillation dynamics near the threshold. At the theoretically predicted threshold $|J|=|J_{\mathrm{b}}|$, the TVM model’s time trace of $\dot{\Phi}_{\mathrm{c}}$ exhibits a characteristic comb-like shape, signifying a strongly non-harmonic oscillation (see the lower panel in Fig. 15(b)). This non-harmonicity occurs because the particle’s energy is extremely close to the potential minimum of the state s Chen et al. (2023b), forcing the particle to traverse the potential valley at a velocity much faster than the velocity at the potential peak.
In contrast, at the macrospin’s higher threshold $|J|=|J^{\prime}_{\mathrm{b}}|$, the particle’s energy is sufficiently far from the potential minimum. Therefore, the time traces of the stable $\dot{\Phi}_{\mathrm{c}}$ (or $M_{z}$) are closer to a sinusoidal (harmonic) shape, as illustrated by the lower panel in Fig. 15(c). This indicates that the particle’s velocities when passing through the potential valley and crossing over the potential peak are now comparable.
@@ III Summary and Discussion
Finally, a pivotal extension of our theory is its scalability to massive multi-lattice systems. Inspired by the ’ionic core’ approximation in atomic physics, we suggest a hierarchical Hamiltonian approach that partitions the system into a highly-constrained internal part (locking sub-lattices into rigid units) and an external part governing collective interactions. By analyzing the local Hamiltonian under periodic constraints, we demonstrate that the collective RBP state s remains the most stable configuration. Local fluctuations, analogous to localized wave packets, are systematically suppressed by the synergy of SOTs and damping torques. This mode-selection mechanism acts as a dynamic stabilizer, damping out high-frequency internal RM modes and forcing the massive lattice to collapse into a coherent assembly of CM units, whose long-range evolution can be effectively described by a unified micromagnetic framework.
@@ Appendix D Generalization of the TVM Model from a Single Lattice to Multiple Lattices
In summary, the TVM framework enables a rigorous coarse-graining transition from atomic lattices to micromagnetic scales. By characterizing each exchange-locked lattice as a robust ’CM-block’-functionally equivalent to a single-domain ferromagnetic unit-we can collapse the high-dimensional internal dynamics into a singular collective coordinate. This mapping allows us to treat large-scale magnetic textures (e.g., poly-crystalline grains or domains) as assemblies of discrete CM-blocks, where complex inter-block interactions, such as long-range exchange and Dzyaloshinskii-Moriya interactions (DMI)Yamane et al. (2019); Wu et al. (2024), are governed by their respective CM variables. This approach not only bridges the gap between atomic-scale LLG simulations and universal oscillator models but also provides a versatile analytical foundation for predicting the collective behavior of large-scale, non-collinear, and AFM/FiM sub-THz devices.
Data Availability Statement
The high-resolution movies (approximately 380 MB) supporting the findings of this study, including animations of degenerate rigid body precessional (RBP) states and RM burst processes, are openly available on Figshare at https://doi.org/10.6084/m9.figshare.31153270.

### v4

@@ II.6 Threshold and Critical Current Densities Verified by the TVM/Macrospin Simulations
Furthermore, the model comparison provides insight into the oscillation dynamics near the threshold. At the theoretically predicted threshold $|J|=|J_{\mathrm{b}}|$, the TVM model’s time trace of $\dot{\Phi}_{\mathrm{c}}$ exhibits a characteristic comb-like shape, signifying a strongly non-harmonic oscillation (see the lower panel in Fig. 15(b)). This non-harmonicity occurs because the particle’s energy is extremely close to the potential maximum of the state s ($E_{\mathrm{c}}=0$) Chen et al. (2023b), forcing the particle to traverse the potential valley at a velocity much faster than the velocity at the potential peak.
In contrast, at the macrospin’s higher threshold $|J|=|J^{\prime}_{\mathrm{b}}|$, the particle’s energy is sufficiently far from the potential maximum. Therefore, the time traces of the stable $\dot{\Phi}_{\mathrm{c}}$ (or $M_{z}$) are closer to a sinusoidal (harmonic) shape, as illustrated by the lower panel in Fig. 15(c). This indicates that the particle’s velocities when passing through the potential valley and crossing over the potential peak are now comparable.
@@ III Summary and Discussion
Finally, a pivotal extension of our theory lies in its multi-scale scalability to massive multi-lattice systems. We propose a hierarchical analytical framework that partitions the system’s complexity into two self-consistent layers. At the fundamental layer, by applying periodic constraints independently to each sub-lattice species, we demonstrate that even when the $120^{\circ}$ rigid-body configuration is distorted by unconventional SOT polarizations or the absence of anisotropy ($H_{\mathrm{u}i}=0$), the system maintains a robust sub-lattice periodic mapping. This allows each micromagnetic cell to be rigorously represented as a composite of three inter-penetrating CM-blocks. At the macroscopic layer, localized fluctuations, analogous to wave packets, are systematically suppressed by the synergic influence of SOTs and damping. This mode-selection mechanism effectively neutralizes high-frequency relative motion (RM) modes, converging the massive lattice into a coherent assembly of CM units. Crucially, as the Poisson Bracket formalism is intrinsically scale-invariant, this framework enables a seamless transition where long-range textures and collective behaviors can be governed by a unified micromagnetic coordination, providing a versatile foundation for the design of large-scale, non-collinear sub-THz devices.
Supplementary Material
See the supplementary material for high-fidelity animations (approximately 380 MB) illustrating the degenerate rigid-body precessional (RBP) states and relative motion (RM) burst processes. These videos provide a direct visualization of the sub-lattice trajectories and the self-resonance mechanism discussed in the Terminal Velocity Motion (TVM) framework.
Data Availability Statement
The high-resolution movies supporting the findings of this study are openly available on Figshare at https://doi.org/10.6084/m9.figshare.31153270. The numerical datasets that support the findings are available from the corresponding author upon reasonable request.
@@ Appendix D Generalization of the TVM Model from a Single Lattice to Multiple Lattices
To self-consistently extend the TVM framework to the micromagnetic scale—particularly in scenarios where SOT excitation may deviate from the standard 120 configuration (state $\mathbf{s}$) due to the absence of uniaxial anisotropy—we adopt a refined coarse-graining strategy. Instead of directly reducing the entire three-spin lattice into a single collective coordinate as in the preceding sections, we preserve the internal degrees of freedom by maintaining periodic boundary conditions (PBC) for each individual sub-lattice species. Specifically, the spins corresponding to the same sub-lattice index across multiple unit cells are first reduced to their respective center-of-mass (CM) coordinates. Consequently, a single micromagnetic cell is represented as a composite assembly of three distinct sub-lattice CM variables. This formulation ensures that the high-frequency relative motions between sub-lattices, inherently present at the atomic scale, are rigorously preserved within the macroscopic micromagnetic framework.
Furthermore, within a singular micromagnetic cell, the system Hamiltonian incorporates the nearest-neighbor exchange interactions between sub-lattice spins, all of which strictly adhere to the aforementioned periodic constraints. Other long-range interactions, including the Dzyaloshinskii-Moriya interaction (DMI) Yamane et al. (2019); Wu et al. (2024), are significantly weaker than the exchange coupling between adjacent lattices; thus, they are insufficient to disrupt the periodic symmetry and effectively become diluted by the ensemble of spin lattices within the cell. Consequently, these higher-order interactions can be elegantly expressed using the collective CM coordinates of the sub-lattices. This hierarchical mapping provides the flexibility to address macroscopic phenomena, such as dynamics in poly-crystalline grains or magnetic domains. Specifically, if the macroscopic SOT excitation preserves the symmetric 120 state ($\mathbf{s}$), the framework allows for a secondary reduction, wherein the TVM model can be derived using a singular collective coordinate at an even larger scale. This multi-layered analytical approach establishes a robust foundation for modeling complex, large-scale NC-AFM devices.